# ChemLog: Making MSOL Viable for Ontological Classification and Learning

### v1

@@ 1 Introduction
Ontologies of structured objects (such as chemical molecules) can be used for the classification of new unseen objects (and thus, for the extension of the ontology) in two ways: (1) the structure of the objects can be used to train deep learning models [7]. These models directly work on a graph-like structure (e.g., Graph Neural Networks), or on a string representation of structured objects (e.g., Transformers). (2) Alternatively, logical axioms can be used for axiomatising classes of structured objects, the latter being construed as first-order models [12]. Note that generally, these logical axioms are not part of the OWL formalisation of the ontology, but are specifications of single classes expressed in a richer language like first-order logic (FOL) or monadic second order logic (MSOL). MSOL has been termed the “finite automata for graphs”, and (unlike FOL) it can be used to speak about paths in a graph and subgraphs of a graph [4]. When using MSOL, a key challenge is its model checking complexity and scalability to large datasets.
@@
In the second part of our paper, we demonstrate how the results of the logic-base classification can improve the quality of deep learning models that classify new molecules in the ChEBI ontology. We train Transformer models both on the original ontology, as well as on our reclassification of the 14 ChEBI classes according to our formalised definitions. The model trained on our reclassified data improves on the ChEBI-trained model by a significant margin (from a micro F1-score of 85% to 93%). Even on the original ChEBI, the new model achieves a recall of 92% (compared to 81% for a model trained directly on ChEBI). This indicates that our classification stays true to the intentions of the original ChEBI classes, while improving the consistency and completeness of the taxonomy.
@@ 2 Axiomatising Peptides
In order to showcase our approach, we will use the concept of peptides in the ChEBI ontology. We have developed MSOL definitions for 14 ChEBI classes related to peptides. Peptides, as defined in ChEBI, are “Amides derived from two or more amino carboxylic acid molecules (the same or different) by formation of a covalent bond from the carbonyl carbon of one to the nitrogen atom of another with formal loss of water. […]”. Thus, our main task for deciding if a molecule is a peptide or not, is to identify the amino acids from which it has been derived. Regarding amino acids, ChEBI states that “The term [peptide] is usually applied to structures formed from $alpha$-amino acids, but it includes those derived from any amino carboxylic acid.”
(a) Structure of the $alpha$-amino acid L-cysteine. All $\alpha$-amino acids follow this structure, only with changing side chains.
(b) Workflow of our methodology. The MSOL axioms are used to verify their correspondences in FOL, which in turn are used to verify the algorithmic implementation. At the same time, each layer can be used on a different scale.
Amino acids are defined by the presence of two functional groups, amino groups and carboxylic acids. In $alpha$-amino acids, they are directly connected by a carbon atom (the amino group is located at the $alpha$ position relative to the carboxylic acid, cf. Figure 1(a)). Since other amino acids are explicitly included, we consider every substructure of a molecule in which the residues of a carboxylic acid and an amino group are connected via a chain of carbon atoms an amino acid residue. Then, we verify if these residues are connected according to the definition (i.e, by a bond between the carbonyl carbon of one amino acid residue and the nitrogen atom of another). Formally, amino acid residues are defined in MSOL as follows:
$\text{{AAR}}(X)\leftrightarrow(\text{{BuildingBlock}}(X)\wedge\exists a,c_{1},c_{2},c_{3}:(\text{{AminoResidue}}(a)\\ \wedge\text{{CarboxyResidue}}(c_{1},c_{2},c_{3})\wedge a\in X\wedge c_{1}\in X\wedge c_{2}\in X\wedge c_{3}\in X))$ (1)
where the building blocks are the fragments of the molecule that are connected via carbon atoms. For a set of carbon atoms $C$, the notion of a connected subset can be expressed as:
$\text{{CarbonConnected}}(X)\leftrightarrow(X\subseteq C\wedge\forall A,B:(A\neq\emptyset\wedge B\neq\emptyset\wedge A\neq B\wedge A\cup B=X)\\ \rightarrow\exists u\in A,v\in B:\text{{Bond}}(u,v))$ (2)
@@
The final definition of peptides uses a 0-ary predicate. For a peptide, at least two distinct amino acid residues are necessary. The substructure resulting from a connection between amino acid residue has to be an amide.
$\text{{Peptide}}\leftrightarrow(\text{{HasCharge{Neutral}}}\wedge\exists A_{1},A_{2},b_{1},b_{2},b_{3}:(\text{{AAR}}(A_{1})\\ \wedge\text{{AAR}}(A_{2})\wedge\neg\text{{Overlap}}(A_{1},A_{2})\wedge\text{{AmideBond}}(b_{1},b_{2},b_{3})\wedge b_{1}\in A_{1}\wedge b_{3}\in A_{2}))$ (3)
Figure 2: FOL structure for the amino acid L-cysteine (cf. Figure 1(a)).
@@ 3 Logical Methodology
This Section describes how the formal definitions of chemical classes are used as a classification tool that determines if a given molecule is part of a class or not. For this purpose, we have developed the ChemLog framework. Figure 1(b) gives an overview of the framework’s workflow. This workflow consists of three stages. In the first stage, MSOL definitions are used directly. For the second stage, a FOL axiomatisation is developed which can be verified by the MSOL axiomatisation. The third stage is an algorithmic implementation of the MSOL axiomatisation which can in turn be verified by the FOL axiomatisation.
The central idea behind this methodology is that the first stage is a direct implementation of the MSOL definitions. However, it is not performant enough to be used on the scale of, e.g., ChEBI with more than 100,000 molecules (see Section 4). Therefore, the second and third stages are iterative steps towards a faster classification and with that classification of larger datasets. Nevertheless, this does not make the original MSOL stage superfluous. Instead, it is used to validate the correctness of the translation to FOL and the algorithmic stage. While we do not give a formal proof of equivalence between the stages, this allows us to compare the classification results between stages 1, 2 and 3, making the translation’s correctness sufficiently plausible.
@@ 3.1 Reasoning in MSOL
For reasoning in MSOL directly, we use two approaches. The first relies on the MSOL model finder MONA [9]. For a given molecule and chemical class, both the structure of the molecule and the class definition are encoded in MONA’s MSOL syntax. For instance, to check if L-cysteine (cf. Figure 1(a)) is an amino acid residue, this includes among others the following lines:
    var2 Atom, N, A;
@@
The second approach encodes the molecule and class definition into a QBF satisfiability problem. This is possible since the size of the domain is limited to the number of atoms. This allows us to represent extensions of predicates as a conjunction of propositional variables, one for each individual (or for each tuple of individuals, for arity >1). Quantification over first-order variables is represented as conjunction or disjunction, and quantification over second-order variables via a series of quantifications over QBF variables. The above example looks like this in QBF:
$\neg O_{1}\wedge\neg O_{2}\wedge\neg O_{3}\wedge\neg O_{4}\wedge O_{5}\wedge O_{6}\wedge\exists x_{0},x_{1},\ldots,x_{6}.(\ldots\wedge\bigvee_{a=0}^{6}\bigvee_{c0=0}^{6}\bigvee_{c1=0}^{6}\bigvee_{c2=0}^{6}\ldots)$ (4)
The $Atom$ predicate limiting the domain size is now implicit. $O_{i}$ states that atom $i$ is an oxygen atom. $x_{i}$ encodes that atom $i$ is in the second-order variable $X$. Note that $\exists A,C$ is not translated the same way as $\exists X$. Instead, since the represent the functional groups amino residue and carboxy residue which have a known, fixed size (1 atom and 3 atoms), $\exists A,C$ is treated like $\exists a,c_{0},c_{1},c_{2}$, adding a disjunction for each first-order variable. The resulting QBF problems are converted to CNF using a Tseitin transformation [14], preprocessed with Bloqqer [1] and solved by DepQBF [13]. In our implementation, MONA and QBF problems (in the QDIMACS format) are automatically generated from an abstract MSOL representation.
@@ 3.2 From MSOL definitions to FOL definitions
Amino acids, the building blocks of peptides, are defined by two functional groups, an amino group and a carboxylic acid, that are connected via a chain of carbon atoms (cf. Section 2). Crucially, this chain can be arbitrarily long. Such a chain cannot be formalised in FOL. Therefore, when defining amino acids, we can only make statements about amino acids in terms of the carbon-connected “building blocks”, not define the connection relation itself. In contrast to MSOL, in FOL the BuildingBlock (cf. Eq. 1) is not logically defined. Instead, it is a primitive predicate whose extension is computed algorithmically and added to a FOL model (see Section 3.3). The translation from MSOL to FOL is performed automatically in ChemLog, a function for calculating the extension of BuildingBlock has been programmed manually.
@@ 3.3 From MSOL model finding to FOL model checking
For applying the FOL axiomatisation, we could use a standard FOL reasoner and encode the FOL structure with axioms, similar to MONA for MSOL. However, for performance reasons, we use FOL structures (like the example in Figure 2) for a model checking problem. On a limited domain (a single molecule), we verify if a set of properties is fulfilled in a single finite model. As discussed in Section 3.2, the building blocks for amino acids have to be calculated algorithmically. Thus, the structure from Figure 2 is augmented with the carbon-connected building blocks. The building blocks are connected to the atoms they include via a binary predicate. To verify that the structure’s augmented part, i.e., the extension of BuildingBlock, corresponds to the MSOL axiomatisation, MONA model finding can be used either directly on the elements of the extension or for the final classification.
@@ 3.5 From FOL to an Algorithmic Implementation
While FOL model checking is more efficient than MSOL theorem proving, it might still be insufficient for large datasets. For the FOL axiomatisation, the extension of a predicate, BuildingBlock, has already been calculated algorithmically. One could now identify other predicates that slow down model checking. In the extreme case, which will be used in the evaluation, the whole classification pipeline is algorithmic. During the development process, the FOL axiomatisation has been used as a verification tool for the implementation of certain predicates and for the classification as a whole. Overall, this gives us three different strategies (including two different implementations for the MSOL strategy) for molecule classification: Fully reasoning-based, partially algorithmic and fully algorithmic. The purpose of this is to provide validation tools (one strategy or implementation can validate another) and to provide scalability, as different strategies provide access to different reasoning tools or, in the case of the algorithmic implementation, do not rely on reasoning tools at all.
@@ 4.1 Layer Alignment
For this dataset, all 4 methods gave the same results (excluding failures). The failure rates and times per molecule for this evaluation are shown in Table 1. The MONA fails to classify more than half of the molecules. This is caused by constraints of the MONA reasoner, which terminates if the internal graph reaches a certain size or a memory limit is exceeded. For the FOL model checker, we have applied a 30 second timeout, leading to failures for 4 of the molecules.
@@ 5 Enhancing Performance of Deep Learning Models
While ChemLog performs well on peptides, this covers only a fraction of ChEBI. A classification using all ChEBI classes therefore has to rely on machine learning models. That said, we here consider the restriction to classification into peptide classes, keeping in mind that an extension to all of ChEBI is much more easily possible than for ChemLog. We interpret this classification as a multi-label machine learning task, using SMILES, a string representation of molecule graphs. To construct the data set, all SMILES-annotated ChEBI classes were taken as instances and the 14 peptide-related classes were used as labels. We are interested in comparing two different ways of assigning labels to instances: The first uses the subsumption relations provided by ChEBI, the second uses our ChemLog methodology.
@@ 8 Conclusion
In this paper we introduced ChemLog, a methodology that harnesses monadic second-order formalisations for ontology classification via FOL model checking and algorithmic implementations. In a case study on the ChEBI ontology, we have developed MSOL definitions for 14 peptide-related classes. For this example we showed how MSOL definitions may be directly translated to FOL (and an extended FOL structure) as well as be implemented algorithmically. The evaluation on the ChEBI ontology has shown that the three stages of our approach can be used to verify each other and provide an efficient debugging mechanism. Using PubChem, we have demonstrated that ChemLog can be used to scale MSOL definitions up to datasets with more than 100 million instances. Finally, we use the results of the logic-based classification to improve the quality of deep learning models. In future work, we will automate the transition from logically defined predicates to algorithmic definitions, using the experience gathered on peptides. This will also allow us to scale up our approach towards other classes in ChEBI as well as ontologies of structured entities in different domains.

### v2

@@ 1 Introduction
The ontological classification of chemical molecules is important for the prediction of chemical properties of molecules, as well as for the discovery of new chemical substances enjoying certain properties. Ontologies of structured objects (such as chemical molecules) can be used for the classification of new unseen objects (and thus, for the extension of the ontology) in two ways. (1) The structure of the objects can be used to train deep learning models [7]. These models directly work on a graph-like structure (e.g., Graph Neural Networks), or on a string representation of structured objects (e.g., Transformers). (2) Alternatively, logical axioms can be used for axiomatising classes of structured objects, the latter being construed as first-order models [12]. Note that generally, these logical axioms are not part of the OWL formalisation of the ontology, but are specifications of single classes. The reason is that they work on a different level: the domain of discourse is no longer the set of all molecules in the ontology, but rather the set of all atoms in a molecule. These specifications are expressed in a richer language like first-order logic (FOL) or monadic second order logic (MSOL). MSOL has been termed the “finite automata for graphs”, and (unlike FOL) it can be used to speak about paths in a graph and subgraphs of a graph [4]. Many important chemical classes like peptides are defined through the presences of certain subgraphs and hence cannot be axiomatised in FOL, leave alone in in OWL. When using MSOL, a key challenge is its model checking complexity and scalability to large datasets.
@@
In the second part of our paper, we demonstrate how the results of the logic-base classification can improve the quality of deep learning models that classify new molecules in the ChEBI ontology. We train Transformer models (ELECTRA [3]) both on the original ontology, as well as on our reclassification of the 14 ChEBI classes according to our formalised definitions. The model trained on our reclassified data improves on the ChEBI-trained model by a significant margin (from a micro F1-score of 85% to 93%). Even on the original ChEBI, the new model achieves a recall of 92% (compared to 81% for a model trained directly on ChEBI). This indicates that our classification stays true to the intentions of the original ChEBI classes, while improving the consistency and completeness of the taxonomy.
@@ 2 Axiomatising Peptides
In order to showcase our approach, we will use the concept of peptides in the ChEBI ontology. We have developed MSOL definitions for 14 ChEBI classes related to peptides. These classes have been selected because peptides play a central role in ChEBI (about 7.4% of ChEBI classes are subclasses of peptide). However, many molecules in this section of ChEBI are underspecified. Often, they belong to a specific subclass of peptide, e.g. tripeptide, but are only classified as peptides. Peptides, as defined in ChEBI, are “Amides derived from two or more amino carboxylic acid molecules (the same or different) by formation of a covalent bond from the carbonyl carbon of one to the nitrogen atom of another with formal loss of water. […]”. Thus, our main task for deciding if a molecule is a peptide or not, is to identify the amino acids from which it has been derived. Regarding amino acids, ChEBI states that “The term [peptide] is usually applied to structures formed from alphaalpha-amino acids, but it includes those derived from any amino carboxylic acid.”
Figure 1: Left: Structure of the alphaalpha-amino acid L-cysteine. All $\alpha$-amino acids follow this structure, only with changing side chains. Right: Workflow of our methodology. The MSOL axioms are used to verify their correspondences in FOL, which in turn are used to verify the algorithmic implementation. At the same time, each layer can be used on a different scale.
Amino acids are defined by the presence of two functional groups, amino groups (predicate name: AminoResidue) and carboxylic acids (CarboxyResidue). In alphaalpha-amino acids, they are directly connected by a carbon atom (the amino group is located at the alphaalpha position relative to the carboxylic acid, cf. Figure 1, left). Since other amino acids are explicitly included, we consider every substructure of a molecule in which the residues of a carboxylic acid and an amino group are connected via a chain of carbon atoms an amino acid residue. The predicate BuildingBlock is used to describe fragments of a molecule that are connected via a chain of carbon atoms. Formally, amino acid residues (AAR) are defined in MSOL as follows:
$\begin{split}\text{{AAR}}(X)\leftrightarrow&(\text{{BuildingBlock}}(X)\wedge\exists a,c_{1},c_{2},c_{3}:(\text{{AminoResidue}}(a)\\ &\wedge\text{{CarboxyResidue}}(c_{1},c_{2},c_{3})\wedge a\in X\wedge c_{1}\in X\wedge c_{2}\in X\wedge c_{3}\in X))\end{split}$ (1)
For a set of carbon atoms $C$, the notion of a connected subset can be expressed as:
$\begin{split}\text{{CarbonConnected}}(X)\leftrightarrow(X\subseteq C\wedge\forall A,B:&(A\neq\emptyset\wedge B\neq\emptyset\wedge A\neq B\wedge A\cup B=X)\\ &\rightarrow\exists u\in A,v\in B:\text{{Bond}}(u,v))\end{split}$ (2)
@@
The final definition of peptides uses a 0-ary predicate (Peptide). For a peptide, two distinct (i.e., not overlapping) amino acid residues have to be connected by an amide bond (AmideBond).
$\begin{split}\text{{Peptide}}\leftrightarrow&(\text{{HasCharge{Neutral}}}\wedge\exists A_{1},A_{2},b_{1},b_{2},b_{3}:(\text{{AAR}}(A_{1})\wedge\text{{AAR}}(A_{2})\\ &\wedge\neg\text{{Overlap}}(A_{1},A_{2})\wedge\text{{AmideBond}}(b_{1},b_{2},b_{3})\wedge b_{1}\in A_{1}\wedge b_{3}\in A_{2}))\end{split}$ (3)
Figure 2: FOL structure for the amino acid L-cysteine (cf. Figure 1).
@@ 3 Logical Methodology
This Section describes how the formal definitions of chemical classes are used as a classification tool that determines if a given molecule is part of a class or not. For this purpose, we have developed the ChemLog framework. Figure 1 (right) gives an overview of the framework’s workflow. This workflow consists of three stages. In the first stage, MSOL definitions are used directly. For the second stage, a FOL axiomatisation is developed which can be verified by the MSOL axiomatisation. The third stage is an algorithmic implementation of the MSOL axiomatisation which can in turn be verified by the FOL axiomatisation.
The central idea behind this methodology is that the first stage is a direct implementation of the MSOL definitions. However, it is not performant enough to be used on the scale of, e.g., ChEBI with more than 100,000 molecules (see Section 4). Therefore, the second and third stages are iterative steps towards a faster classification and with that classification of larger datasets. Nevertheless, this does not make the original MSOL stage superfluous. Instead, it is used to validate the correctness of the FOL axiomatisation and of the algorithmic stage. While we do not give a formal proof of equivalence between the stages, this allows us to compare the classification results between stages 1, 2 and 3, making the translation’s correctness sufficiently plausible.
@@ 3.1 Reasoning in MSOL
For reasoning in MSOL directly, we use two approaches. The first relies on the MSOL model finder MONA [9]. For a given molecule and chemical class, both the structure of the molecule and the class definition are encoded in MONA’s MSOL syntax. For instance, to check if L-cysteine (cf. Figure 1) is an amino acid residue, this includes among others the following lines:
    var2 Atom, O, A;
@@
The second approach encodes the molecule and class definition into a Quantified Boolean Formula (QBF) satisfiability problem. This is possible since the size of the domain is limited to the number of atoms. This allows us to represent extensions of predicates as a conjunction of propositional variables, one for each individual (or for each tuple of individuals, for arity >1). Quantification over first-order variables is represented as conjunction or disjunction, and quantification over second-order variables via a series of quantifications over QBF variables. The above example looks like this in QBF:
$\begin{split}&\text{{Atom}}_{0}\wedge\text{{Atom}}_{1}\wedge\text{{Atom}}_{2}\wedge\text{{Atom}}_{3}\wedge\text{{Atom}}_{4}\wedge\text{{Atom}}_{5}\wedge\text{{Atom}}_{6}\wedge\neg O_{0}\wedge\neg O_{1}\wedge\neg O_{2}\\ &\wedge\neg O_{3}\wedge\neg O_{4}\wedge O_{5}\wedge O_{6}\wedge\exists x_{0},x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}:(\ldots\wedge\bigvee_{a=0}^{6}\bigvee_{c0=0}^{6}\bigvee_{c1=0}^{6}\bigvee_{c2=0}^{6}(\ldots))\end{split}$ (4)
$Atom_{i}$ states that the $i$-th element in the domain is an atom. $O_{i}$ states that atom $i$ is an oxygen atom. The last term of the disjunction encodes the definition of amino acid residues (Eq. 1). $x_{i}$ encodes the presence or absence of atom $i$ in the second-order variable $X$. $\exists a$ is represented with the disjunction $\bigvee_{a=0}^{6}$. The resulting QBF problems are converted to CNF using a Tseitin transformation [14], preprocessed with Bloqqer [1] and solved by DepQBF [13]. In our implementation, MONA and QBF problems (in the QDIMACS format) are automatically generated from an abstract MSOL representation.
@@ 3.2 From MSOL Definitions to FOL Definitions
Amino acids can have an arbitrarily long chain of carbon atoms connecting the amino group and the carboxylic acid (cf. Section 2). Such a chain cannot be formalised in FOL. Therefore, when defining amino acids, we can only make statements about amino acids in terms of the carbon-connected “building blocks”, not define the connection relation itself. In contrast to MSOL, in FOL the BuildingBlock (cf. Eq. 1) is not logically defined. Instead, it is a primitive predicate whose extension is computed algorithmically and added to a FOL model (see Section 3.3). The translation from MSOL to FOL is performed automatically in ChemLog. The only predicate that is not automatically translated is BuildingBlock, the extension of which is calculated by a manually implemented Python function.
@@ 3.3 From MSOL Model Finding to FOL Model Checking
For FOL, model checkers are available and easily implemented. Hence we can directly use the original problem formulation and check whether the FOL axiomatisation of a given class holds in a FOL structure that represents a molecule (like the example in Figure 2). As discussed in Section 3.2, the building blocks for amino acids have to be calculated algorithmically. Thus, the structure from Figure 2 is augmented with the carbon-connected building blocks. The building blocks are connected to the atoms they include via a binary predicate. To verify that the structure’s augmented part, i.e., the extension of BuildingBlock, corresponds to the MSOL axiomatisation, MONA model finding can be used either directly on the elements of the extension or for the final classification.
@@ 3.5 From FOL to an Algorithmic Implementation
While FOL model checking is more efficient than MSOL theorem proving, it might still be insufficient for large datasets. Therefore, we have implemented an automatic translation from FOL to Python. Each predicate corresponds to a Python function which accepts an RDKit  [^1] https://www.rdkit.org/ molecule object (mol), the second-order objects that have already been calculated algorithmically (i.e., the extension of BuildingBlock, cf. Section 3.2, so_elements) and the arguments of the FOL predicate. For example, for amino acid residues, we get the following Python function:
def AAR(mol: Chem.Mol, so_elements, x):
    return any((x in so_elements and AminoResidue(mol, so_elements, x0)
        and x0.GetIdx() in x and CarboxyResidue(mol, so_elements, x1, x2,
        x3) and x1.GetIdx() in x) for x0 in mol.GetAtoms() for x1 in
        mol.GetAtoms() for x2 in mol.GetAtoms() for x3 in mol.GetAtoms())
The final algorithmic implementation has been further optimised for efficiency. Overall, this gives us three different strategies (including two different implementations for the MSOL strategy) for molecule classification: Fully reasoning-based, partially algorithmic and fully algorithmic. The purpose of this is to provide validation tools (one strategy or implementation can validate another) and to provide scalability, as different strategies provide access to different reasoning tools or, in the case of the algorithmic implementation, do not rely on reasoning tools at all. The implementation for ChemLog is available on GitHub  [^2] https://github.com/sfluegel05/chemlog-peptides. A web front-end for molecule classification can also be accessed as part of Chebifier [7]  [^3] https://chebifier.hastingslab.org/. Select “ChemLog Peptides” to get only results from ChemLog as discussed in this work.
@@ 4.1 Layer Alignment
For this dataset, all 4 methods gave the same results (excluding failures). The failure rates and times per molecule for this evaluation are shown in Table 1. The MONA fails to classify more than half of the molecules. This is caused by non-configurable constraints of the MONA reasoner, which terminates if the internal graph reaches a certain size or a memory limit is exceeded. For the FOL model checker, we have applied a 30 second timeout, leading to failures for 4 of the molecules.
@@ 5 Enhancing Performance of Deep Learning Models
Recall that our task is the classification of (new) molecules into ChEBI classes. Since these classes can overlap, this is a multi-label classification problem. While ChemLog performs well on peptides, this covers only a fraction of ChEBI. Classification based on the ChemLog methodology requires the design of a large amount of logical axioms, which currenly does not scale to the whole of ChEBI. Therefore, we also have to rely on machine learning models. That said, we here consider the restriction to classification into peptide classes, keeping in mind that an extension to all of ChEBI is much more easily possible for machine learning than for ChemLog. We interpret this classification as a multi-label machine learning task, using SMILES, a string representation of molecule graphs, as inputs. To construct the data set, all SMILES-annotated ChEBI classes were taken as instances and the 14 peptide-related classes were used as labels. We are interested in comparing two different ways of assigning labels to instances: The first uses the subsumption relations provided by ChEBI, the second uses our ChemLog methodology.
@@ 8 Conclusion
In this paper we introduced ChemLog, a methodology that harnesses monadic second-order formalisations for ontology classification via FOL model checking and algorithmic implementations. In a case study on the ChEBI ontology, we have developed MSOL definitions for 14 peptide-related classes. For this example we showed how MSOL definitions may be directly translated to FOL (and an extended FOL structure) as well as be implemented algorithmically. The evaluation on the ChEBI ontology has shown that the three stages of our approach can be used to verify each other and provide an efficient debugging mechanism. Using PubChem, we have demonstrated that ChemLog can be used to scale MSOL definitions up to datasets with more than 100 million instances. Finally, we use the results of the logic-based classification to improve the quality of deep learning models. In future work, we will improve the transition from logically defined predicates to algorithmic definitions, using the experience gathered on peptides. This will include automating the translation of second-order predicates and improving the efficiency of the FOL definitions and Python functions resulting from the translation. The general methodology employed in this paper does not make use of specific properties of peptides. Hence, we expect that we can scale up our approach towards other classes in ChEBI as well as ontologies of structured entities in different domains.